OpenStudy (anonymous):

Tom and sam sit down to eat a pie. Tom eats 2/3 of the pie, same eats 1/4 of the pie. What fraction of the pie is left?

OpenStudy (anonymous):

1 - (2/3 + 1/4) = 1 - (8/12 + 3/12) = 1 - (11/12) = 1/12

OpenStudy (anonymous):

The key to this is converting fractions. Specifically, converting 2/3 to 8/12 and 1/4 to 3/12. This is done by finding the LCM of 3 and 4.